The mean and variance of defective items is 0.72 and 0.6876. Find the probability of getting 12 non-defective items.


1
Expert's answer
2022-03-28T05:48:12-0400

"Z=\frac{x-\mu}{\sigma}=\frac{12-0.72}{\sqrt{0.6876}}=13.6"

P(Z=13.6)=0